1. Use the definition of the limit  to show that the sequence (1 + (−1)^n) is divergent.
1
Expert's answer
2021-05-07T12:29:49-0400

"\begin{array}{l}
\sum_{n=0}^{\infty}\left(1+(-1)^{n}\right) \rightarrow \text { divergent } \\
\text { Rule: We have } \sum_{k=0}^{\infty} a_{k} \text { , } \\
\text { if } \lim _{k \rightarrow \infty} a_{k}=0 \text { , the suries diverges } \\
\lim _{n \rightarrow \infty}\left(1+(-1)^{n}\right)=\lim _{n \rightarrow \infty} 1+\lim _{n \rightarrow \infty}(-1)^{n}= \\
=1+1=2 \\
n=2 k, k \in N \\
\text { the series diverges }
\end{array}"